In 2010 a survey of 600 Twilight fans found 378 said that they would prefer to spend eternity with Jacob (over Edward). You work for Twilight and wish to estimate, with 99% confidence, the proportion of Twilight fans who favor eternity with Jacob (over Edward) in 2020. Your estimate must be accurate within 3% of the true value. Find the minimum sample size needed.
Answer #1

The sample proportion here is computed as:

p = x/n = 378/600 = 0.63

From standard normal tables, we have:
P( -2.576 < Z < 2.576) = 0.99

Therefore the margin of error here is computed as:

Putting all the given values, we get here:

Therefore 1719 is the required sample size here.